OpenStudy (anonymous):

What is the arc length when Θ = 4 pi over 7 and the radius is 5 cm?

OpenStudy (anonymous):

@HELP!!!!

OpenStudy (anonymous):

\[\frac{arc}{7}=\frac{4\pi}{7}\] \[arc=4\pi\]

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

20 pi over 7cm 4 pi over 7cm 2 pi over 7cm 24 pi over 7cm

OpenStudy (anonymous):

@ HELP!!!! can you help me with this one?

OpenStudy (anonymous):

The function f(t) = 5 cos(pi over 6t) + 7 represents the tide in Stanley Sea. It has a maximum of 12 feet when time (t) is 0 and a minimum of 2 feet. The sea repeats this cycle every 12 hours. After four hours, how high is the tide?

OpenStudy (anonymous):

@HELP!!!!

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@ganeshie8